OpenStudy (anonymous):

how do you solve this: log(x+1) = log(2)- log(4x-1)

OpenStudy (alfie):

\[\log(x-1) = \log \frac{2}{4x-1}\] x-1 = 2/(4x-1) Solve for x.

OpenStudy (mathteacher1729):

Probably add log ( 4x - 1) to both sides to obtain: log(x+1) + log(4x-1) = log(2) now use the product rule for logs to obtain: log( (x+1)(4x-1) ) = log(2) Assuming the bases of the logs are the same of course. Now you just solve the quadratic equation: (x+1)(4x-1) = 2 Which you can do by expanding, collecting terms, and factoring again. :)

OpenStudy (alfie):

sorry I misread the plus with the minus up there, but it still works ;) bye o/